OpenStudy (anonymous):

12*​x^​5-​x^​3+​6*​x/​(4*​x^​3

OpenStudy (anonymous):

\[12​x^5−​x^3+​\frac{ 6x }{ 4x​^3} or \frac{ 12​x^5−​x^3 +6x }{ 4x​^3 }\]

OpenStudy (anonymous):

and what do you want to know about it? derivatives zero points graph?

OpenStudy (anonymous):

12x^5-x^3+6x divided by 4x^3 is the problem and i have no i dea what the firsts step in solving it is

OpenStudy (anonymous):

xo the second? \[\frac{ 12​x^5−​x^3+6x }{ 4x​^3 }\] what does it equal? zero?

OpenStudy (anonymous):

what do you want to solve it for, what does it equate? this is just a function

OpenStudy (anonymous):

yes, i want to know what it equals.

OpenStudy (anonymous):

That's what you need to give, Or is there some domain for x given?

OpenStudy (anonymous):

that is some values of x? Is that given?

OpenStudy (anonymous):

what? no i just need to simplify the problem as much as possible.

OpenStudy (anonymous):

@abster i apologize if i am confusing you, but the values of x are not given, you just have to simplify to get the answer

OpenStudy (anonymous):

OK, just a sec

OpenStudy (anonymous):

\[\frac{ 12​x^5-​x^33+​6​x }{ 4x^3 } =\frac{ 12​x^​4-​x^​2+​6x }{ 4x^2 }\]

OpenStudy (anonymous):

is the first step (scratch one x

OpenStudy (anonymous):

you wouldnt divide 12 and 4?

OpenStudy (anonymous):

sorry I made a mistake \[\frac{ 12​x^5-​x^33+​6​x }{ 4x^3 } =\frac{ 12​x^​4-​x^​2+​6 }{ 4x^2 }\\]]

OpenStudy (anonymous):

\[\frac{ 12​x^5-​x^33+​6​x }{ 4x^3 } =\frac{ 12​x^​4-​x^​2+​6 }{ 4x^2 }\]

OpenStudy (anonymous):

okay so how did you get x^2+6?

OpenStudy (anonymous):

that stupid equation editer is not behaving very well \[\frac{ 12​x^5-​x^3+​6​x }{ 4x^3 } =\frac{ 12​x^4-​x^2+​6 }{ 4x^2 } =3x^2 -\frac{ 1 }{4 } +\frac{ 1.5 }{x^2 }\]

OpenStudy (anonymous):

It depends what you consider simple

OpenStudy (anonymous):

Oh, and it is only if x is unequal to zero, (there the function does not exist

OpenStudy (anonymous):

oh okay, i think i understand, thanks so much!

OpenStudy (anonymous):

if you would want to know the zeropoints, you might be after something like (ax² + b)(cx² + d)/4x² = (acx⁴ +(bc + ad)x² + bd)/4x² had you have to find a, b, c, and d such that ac = 12 bc + ad = -1 bd = 6 but I don't see an easy solution to that, so I don't think that is what they are after